# Is neural semantic parsing good at ellipsis resolution, or isn’t it?

### v2

@@ Abstract
Neural semantic parsers have shown good overall performance for a variety of linguistic phenomena, reaching semantic matching scores of more than 90%. But how do such parsers perform on strongly context-sensitive phenomena, where large pieces of semantic information need to be duplicated to form a meaningful semantic representation? A case in point is English verb phrase ellipsis, a construct where entire verb phrases can be abbreviated by a single auxiliary verb. Are the otherwise known as powerful semantic parsers able to deal with ellipsis or aren’t they? We constructed a corpus of 120 cases of ellipsis with their fully resolved meaning representation and used this as a challenge set for a large battery of neural semantic parsers. Although these parsers performed very well on the standard test set, they failed in the instances with ellipsis. Data augmentation helped improve the parsing results. The reason for the difficulty of parsing elided phrases is not that copying semantic material is hard, but that usually occur in linguistically complicated contexts causing most of the parsing errors.
@@ 1 Introduction
Semantic parsing is the task of providing a formal meaning representation for an input sentence of a natural language such as English, Dutch, or Italian. Semantic parsing is crucial for applications that require the precise translation of unstructured data (i.e., text and images) into structured data (e.g., databases and robot commands). Currently, the most promising approaches to semantic parsing are based on neural models Bai et al. (2022); Wang et al. (2023); Zhang et al. (2024b, 2025) trained or fine-tuned on large semantically annotated corpora Banarescu et al. (2013); Abzianidze et al. (2017a), reaching high performance with F scores greater than 90%. Little is known about the ability of neural semantic parsers to cope with ellipsis, a linguistic construction in which elements are omitted and are supplied by the discourse context. In this paper, we will study how neural semantic parsers deal with Verb Phrases Ellipsis (VPE) in English. An example of a VPE is shown in (1) together with its fullly expressed surface interpretation in (2).
@@ 3 Data
The PMB Abzianidze et al., 2017b is a multilingual corpus enriched with semantic annotations, covering a wide range of linguistic phenomena. It contains a substantial set of parallel texts, each paired with a formal meaning representation known as a Discourse Representation Structure (DRS) based on Discourse Representation Theory (DRT, Kamp and Reyle, 1993). While DRSs are typically presented in a human-readable box format, a clause-based linear representation was introduced by van Noord et al. (2018) to enable their use in sequence-based models. More recently, Bos (2023) proposed Sequence Box Notation (SBN), a simplified, variable-free version of DRS aimed at further facilitating sequence processing. In this paper we use SBN as meaning representation format (see Figure 1).
@@
As occurrences of VPE are relatively rare Bos and Spenader (2011), it is rather challenging to yield a reasonably sized corpus. A total of 120 cases were identified in the PMB and their corresponding meaning representations manually corrected. Slightly more than half of the cases (71) contained some kind of negation in the elided construction (e.g., ”and neither am I”, ”Greenland is not”, ”but she didn’t”). Half of the ellipsis are escorted by the auxiliary verb to do, a third by to be, and the remaining cases are formed by other auxiliary verbs, the infinitival particle to or instances of gapping. An annotated example taken from the corpus is shown in Figure 1, where the elliptical phrase ”life does” is semantically interpreted as ”life does end”.
@@ 4 Experimental Setup
We evaluate model performance using two metrics: Smatch and Ill-Formed Rate (IFR). Smatch Cai and Knight (2013); Zhang et al. (2024b) measures the similarity between the predicted and reference semantic graphs by converting each graph into a set of triples and computing the optimal variable mapping via a hill-climbing algorithm. Precision (P), recall (R), and F1 score are calculated as follows:
@@ 5 Results and Analysis
mBart-Large Standard 79.50 3.95 65.90 29.17
Aug300 81.40 4.00 72.90 23.33
Aug600 81.00 3.60 73.10 20.83
mT5-Large Standard 78.61 8.20 65.38 25.83
Aug300 80.50 6.80 70.20 20.83
Aug600 80.00 6.20 70.50 20.00
byT5-Large Standard 87.40 5.73 61.22 23.33
Aug300 88.50 4.50 68.00 18.33
Aug600 88.90 4.00 67.50 18.33
Qwen2.5-7B Standard 91.19 2.34 72.09 13.33
Aug300 91.35 2.17 80.31 5.83
Aug600 91.57 2.09 79.64 8.33
Ministral-8B Standard 93.06 1.67 77.77 9.17
Aug300 93.28 1.25 84.00 2.50
Aug600 93.01 1.59 85.61 2.50
LLaMA3.1-8B Standard 92.18 1.51 78.11 8.33
Aug300 91.61 2.18 83.89 8.33
Aug600 92.08 1.76 84.21 4.17
Gemma2-9B Standard 92.80 1.59 73.09 13.33
Aug300 93.02 1.42 83.52 3.33
Aug600 93.26 1.09 84.20 4.17
@@
Decoder-only architectures consistently outperform encoder-decoder models on the Standard Test set. Ministral-8B achieves the highest performance with a Smatch score of 93.28 following augmentation (compared to 93.06 on the standard training set). Other decoder-only models demonstrate similarly strong performance: Gemma2-9B (92.80), LLaMA3.1-8B (92.18), and Qwen2.5-7B (91.19) all maintain scores above 91. In contrast, encoder-decoder architectures (mBART-Large, mT5-Large, and byT5-Large) achieve lower performance, with standard scores ranging from 78 to 87. This performance disparity likely stems from both architectural differences and parameter scale advantages, where larger decoder-only models may benefit from more stable fine-tuning dynamics and in-context learning capabilities.
@@
All models exhibit substantially degraded performance on VPE120 relative to the Standard Test, confirming the inherent difficulty of parsing elliptical constructions semantically. When trained solely on the standard dataset, models achieve VPE120 Smatch scores between 61.22 and 78.11, accompanied by markedly increased IFR (e.g., 29.17% for mBART-Large and 25.83% for mT5-Large), indicating frequent generation of malformed outputs.
VPE-specific data augmentation yields substantial improvements across all architectures. Ministral-8B achieves the highest score of 85.61 with Aug600, closely followed by Gemma2-9B (84.20) and LLaMA3.1-8B (84.21). These top-performing models also demonstrate the most significant IFR reductions (e.g., Ministral-8B: 9.17% to 2.50%). Encoder-decoder models also benefit from augmentation: mBART-Large improves from 65.90 (standard) to 73.10 (Aug600), while mT5-Large advances from 65.38 to 70.50. Notably, byT5-Large shows improvement from 61.22 to 68.00 with Aug300.
@@
The previous section showed that sentences with ellipsis are a lot harder to parse for the neural semantic models. But why is this is the case? Is this because they are a bad at copying semantic information, or is it something else? In order to answer we examined the output of the best performing model and manually inspected the results.
Surprisingly, what we thought would be hard for the models, copying semantic material from the source to the target, was not hard at all. Only in three of the 120 cases did this not happen. Actually, what contributed to the low score was the wrong choice of discourse relation (22% of overall errors), the wrong attachment of a discourse relation (20%), incorrect scope order between tense and negation (16%) incorrect choice of word sense (16%), incorrect choice of thematic role (10%), incorrect choice of concept (10%), and incorrectly resolved anaphora (4%).
@@ 6 Conclusion
Although open-domain semantic parsing achieves good overall performance on the standard test sets, there shortcomings arise at the surface when looking at more complex linguistic phenomena. We demonstrated this by looking specifically at how neural parsing models deal with English VP Ellipsis. Although we observed a drop in performance, the reason for the drop was not the context-sensitive nature of ellipsis, but rather the fact that elliptical phenomena are often surrounded by complex phenomena such as tense, negation, and discourse structure, causing parsing errors.

### v3

@@ Abstract
Neural semantic parsers have shown good overall performance for a variety of linguistic phenomena, reaching semantic matching scores of more than 90%. But how do such parsers perform on strongly context-sensitive phenomena, where large pieces of semantic information need to be duplicated to form a meaningful semantic representation? A case in point is English verb phrase ellipsis, a construct where entire verb phrases can be abbreviated by a single auxiliary verb. Are the otherwise known as powerful semantic parsers able to deal with ellipsis or aren’t they? We constructed a corpus of 120 cases of ellipsis with their fully resolved meaning representation and used this as a challenge set for a large battery of neural semantic parsers. Although these parsers performed very well on the standard test set, they failed in the instances with ellipsis. Data augmentation helped improve the parsing results. The reason for the difficulty of parsing elided phrases is not that copying semantic material is hard, but that they usually occur in linguistically complicated contexts, causing most of the parsing errors.
@@ 1 Introduction
Semantic parsing is the task of providing a formal meaning representation for an input sentence of a natural language such as English, Dutch, or Italian. Semantic parsing is crucial for applications that require the precise translation of unstructured data (i.e., text and images) into structured data (e.g., databases and robot commands). Currently, the most promising approaches to semantic parsing are based on neural models Bai et al. (2022); Wang et al. (2023); Zhang et al. (2024b, 2025) trained or fine-tuned on large semantically annotated corpora Banarescu et al. (2013); Abzianidze et al. (2017a), reaching high performance with F scores greater than 90%. Little is known about the ability of neural semantic parsers to cope with ellipsis, a linguistic construction in which elements are omitted and are supplied by the discourse context. In this paper, we will study how neural semantic parsers deal with Verb Phrase Ellipsis (VPE) in English. An example of a VPE is shown in (1) together with its fully expressed surface interpretation in (2).
@@ 3 Data
The PMB Abzianidze et al., 2017a is a multilingual corpus enriched with semantic annotations, covering a wide range of linguistic phenomena. It contains a substantial set of parallel texts, each paired with a formal meaning representation known as a Discourse Representation Structure (DRS) based on Discourse Representation Theory (DRT, kampreyle:drt). While DRSs are typically presented in a human-readable box format, a clause-based linear representation was introduced by van Noord et al. (2018) to enable their use in sequence-based models. More recently, Bos (2023) proposed Sequence Box Notation (SBN), a simplified, variable-free version of DRS aimed at further facilitating sequence processing. In this paper we use SBN as meaning representation format (see Figure 1).
@@
As occurrences of VPE are relatively rare Bos and Spenader (2011), it is rather challenging to yield a reasonably sized corpus. A total of 120 cases were identified in the PMB and their corresponding meaning representations manually corrected. Slightly more than half of the cases (71) contained some kind of negation in the elided construction (e.g., ”and neither am I”, ”Greenland is not”, ”but she didn’t”). Half of the instances are accompanied by the auxiliary verb to do, a third by to be, and the remaining cases are formed by other auxiliary verbs, the infinitival particle to or instances of gapping. An annotated example taken from the corpus is shown in Figure 1, where the elliptical phrase ”life does” is semantically interpreted as ”life does end”.
@@ 4 Experimental Setup
We evaluate model performance using two metrics: Smatch [^1] We adopt the Smatch++ implementation opitz-2023-smatch, which uses Integer Linear Programming (ILP) instead of the standard hill-climbing approach. and Ill-Formed Rate (IFR). Smatch Cai and Knight (2013); opitz-2023-smatch measures the similarity between the predicted and reference semantic graphs by converting each graph into a set of triples and computing the optimal variable mapping via a hill-climbing algorithm. Precision (P), recall (R), and F1 score are calculated as follows:
@@ 5 Results and Analysis
mBart-Large Standard 83.50 6.95 70.90 33.17
Aug300 85.40 7.00 77.90 27.33
Aug600 85.00 6.60 78.10 24.83
mT5-Large Standard 82.61 11.20 70.38 29.83
Aug300 84.50 9.80 75.20 24.83
Aug600 84.00 9.20 75.50 24.00
ByT5-Large Standard 91.40 8.73 66.22 27.33
Aug300 92.50 7.50 73.00 22.33
Aug600 92.90 7.00 72.50 22.33
Qwen2.5-7B Standard 94.19 5.34 77.09 17.33
Aug300 94.35 5.17 85.31 9.83
Aug600 95.50 5.09 84.64 12.33
Ministral-8B Standard 95.45 4.67 82.77 13.17
Aug300 95.50 4.25 89.00 6.50
Aug600 95.42 4.59 90.61 6.50
LLaMA3.1-8B Standard 95.56 4.51 83.11 12.33
Aug300 95.32 5.18 88.89 12.33
Aug600 95.44 4.76 89.21 8.17
Gemma2-9B Standard 96.31 4.59 78.09 17.33
Aug300 96.46 4.42 88.52 7.33
Aug600 96.59 4.09 89.20 8.17
@@
Decoder-only architectures consistently outperform encoder–decoder models on the Standard Test set. Gemma2-9B achieves the highest performance with a Smatch score of 96.59 following augmentation (compared to 96.31 on the standard training set). Other decoder-only models demonstrate similarly strong performance: LLaMA3.1-8B (95.44), Ministral-8B (95.50), and Qwen2.5-7B (95.50) all maintain scores above 95. In contrast, encoder–decoder architectures (mBART-Large, mT5-Large, and ByT5-Large) achieve lower performance, with standard scores ranging from 82.61 to 91.40. This performance disparity likely stems from both architectural differences and parameter scale advantages, where larger decoder-only models may benefit from more stable fine-tuning dynamics and in-context learning capabilities.
@@
All models exhibit substantially degraded performance on VPE120 relative to the Standard Test, confirming the inherent difficulty of parsing elliptical constructions semantically. When trained solely on the standard dataset, models achieve VPE120 Smatch scores between 66.22 and 83.11, accompanied by markedly increased IFR (e.g., 33.17% for mBART-Large and 29.83% for mT5-Large), indicating frequent generation of malformed outputs.
VPE-specific data augmentation yields substantial improvements across all architectures. Ministral-8B achieves the highest score of 90.61 with Aug600, closely followed by LLaMA3.1-8B (89.21) and Gemma2-9B (89.20). These top-performing models also demonstrate the most significant IFR reductions (e.g., Ministral-8B: 13.17% $\rightarrow$ 6.50%). Encoder–decoder models also benefit from augmentation: mBART-Large improves from 70.90 (standard) to 78.10 (Aug600), while mT5-Large advances from 70.38 to 75.50. Notably, ByT5-Large shows improvement from 66.22 to 73.00 with Aug300.
@@
The previous section showed that sentences with ellipsis are a lot harder to parse for the neural semantic models. But why is this the case? Is this because they are a bad at copying semantic information, or is it something else? In order to answer we examined the output of the best performing model and manually inspected the results.
Surprisingly, what we thought would be hard for the models, copying semantic material from the source to the target, was not hard at all. Only in three of the 120 cases did this not happen. Actually, what contributed to the low score was the wrong choice of discourse relation (22% of overall errors), the wrong attachment of a discourse relation (20%), incorrect scope order between tense and negation (16%), incorrect choice of word sense (16%), incorrect choice of thematic role (10%), incorrect choice of concept (10%), and incorrectly resolved anaphora (4%).
One reason why selecting the correct VP antecedent might have to do with the amount of ambiguity, or lack thereof. For instance, in the VPE example in Figure 1 there is only one potential verb phrase that could serve as antecedent for the elliptical phrase. Closer inspection of the dataset reveals that most (81%) of the texts with VPE are relatively short and provide only one verb phrase that could act as antecedent; only 23 examples provide two or more potential verb phrase antecedents, as in (3).
(3) Ann hoped to succeed, but she didn’t.
Here there are two verb phrases in the context: hope to succeed and succeed. For most of these cases picking the most recent verb phrase usually yields the correct interpretation.
@@ 6 Conclusion
Although open-domain semantic parsing achieves good overall performance on the standard test sets, its shortcomings arise at the surface when looking at more complex linguistic phenomena. We demonstrated this by looking specifically at how neural parsing models deal with cases of English VP Ellipsis. Although we observed a drop in performance, the reason for the drop was not the context-sensitive nature of ellipsis, but rather the fact that elliptical phenomena are often surrounded by complex phenomena such as tense, negation, and discourse structure, causing parsing errors. So, is neural semantic parsing good at ellipsis resolution? Yes, it is!
Acknowledgements
We would like to thank the three anonymous reviewers for their comments. Reviewer 1 pointed out that the developed datasets will be valuable for the community, and indeed the VPE dataset will be made public via the Parallel Meaning Bank data releases. Reviewer 2 wondered how many different VP targets there are for each case of ellipsis, as the complexity for ellipsis resolution ”depends greatly on the number of available VP targets for each VPE”. We added a discussion in this topic in Section 5. Reviewer 3 noted that the question in the paper’s title is not actually answered. This was a correct observation. But now we did explicitly in Section 6. Special thanks go to Juri Opitz, who pointed out that using a hill-climber for evaluation (as we did in the submitted version of this paper) is not optimal. So we recalculated our scores using ILP instead.